A company whose stock is selling for $45 has the following balance sheet: Assets $32,000 Liabilities $10,000 Common stock 6,000 ($6 par; 1,000 shares issued) Additional paid-in 2,000 capital Retained earnings 14,000 ​ a. Construct a new balance sheet showing a 3 for 1 stock split. What is the new price for the stock? ​ b. What would be the balance sheet if the firm paid a 10 percent stock dividend (instead of the stock split)?
Answer #1
Requirement a
New balance sheet after a 3 for 1 Split:
$ $
Assets 32000 Liabilities 10000
Common stock ($2 par; 3000 shares outstanding) 6000
Additional Paid in Capital 2000
Retained Earnings 14000
32000 32000
In stock spilt, there is only change in number of shares and par value
New number of Shares outstanding = 1000 X 3 = 3000 shares
New par value per share = $6/3 = $2 par value per share
Requirement b
New balance sheet after the 10 percent stock dividend:
$ $
Assets 32000 Liabilities 10000
Common stock ($6 par; 1100 shares outstanding) 6600
Additional Paid in Capital 5900
Retained Earnings 9500
32000 32000
Sotck dividend shares = 1000*10% = 100 Shares
New number of common stock shares outstanding = 1000+100 = 1,100 Shares with $6 par value. So common stock is increased by $600 to $6,600
Market value of stock = 100 X 45 = $4,500
Retained earned are reduced from $14,000 to $9,500
Since, Retained earnings are reduced by $4,500 and common stock is increased by $600. So balance $ 3,900 Increase in Additional paid in capital